OpenStudy (anonymous):

Which is the side length of a cube with a volume of 3375 in3? A. B. C. D.

OpenStudy (welshfella):

you need to find the cube root of 3375 grab ur calculator!

OpenStudy (anonymous):

Your answer would be 34, simply because when you use your scientific calc. to find the cubic root of (33750 you'll see that you get this answer

OpenStudy (welshfella):

no its not 34

OpenStudy (briellshaw):

divide you length by four

OpenStudy (briellshaw):

3375 by 4

OpenStudy (jdoe0001):

 \(\bf volume =x\cdot x\cdot x\implies volume =x^3 \\ \quad \\ volume = 3375\implies 3375=x^3\impliedby \textit{solve for }x\)